### v1

# It Takes Two: A Peer-Prediction Solution for Blockchain Verifier’s Dilemma

Abstract
The security of blockchain systems is fundamentally based on the decentralized consensus in which the majority of parties behave honestly, and the process of content verification is essential to maintaining the robustness of blockchain systems. However, the phenomenon that a secure blockchain system with few or no cheaters could not provide sufficient incentive for verifiers to honestly perform the costly verification, referred to as the Verifier’s Dilemma, could severely undermine the fundamental security of blockchain systems. While existing works have attempted to insert deliberate errors to disincentivize lazy verification, the decentralized environment makes it impossible to judge the correctness of verification or detect malicious verifiers directly.
In this paper, we initiate the research that leverages the peer prediction approach towards the design of Bayesian truthful mechanisms for the decentralized verification game among multiple verifiers, incentivizing all verifiers to perform honest verification without access to the ground truth even in the presence of noisy observations in the verification process. With the theoretically guaranteed truthfulness of our mechanism for the verification game, our work provides a framework of verification mechanisms that enhances the security and robustness of the blockchain and potentially other decentralized systems.
1 Introduction
Blockchain, with prevailing examples as Bitcoin [1] and Ethereum [2], is an emerging technology that maintains decentralized consensus via a distributed ledger that utilizes cryptographic techniques to achieve trust and security. To ensure that the blockchain runs correctly, certain contents on the blockchain need to be “proven” and verified by other parties. The “proofs” can either be developed cryptographically, e.g. zero-knowledge proof [3], or in a game-theoretical way, e.g. via a verification game. While the method of zero-knowledge proofs achieves cryptographical security, its high computational overhead limits its application in large-scale tasks, e.g. machine learning [4].
In several previous studies on blockchain verification (e.g. [5]), it is typical that the system only prevents the provers from cheating while assuming that verifiers are honest. However, in a fully decentralized and permissionless blockchain system, this is not necessarily true, especially in verification games in which the incentive-compatibility is not theoretically guaranteed.
When considering game-theoretic ways to incentivize verifiers to verify honestly, a straightforward way is to induce a penalty (“slash”) when a verifier is detected to be dishonest, as in Ethereum [6], opML [7], etc. Nevertheless, a simple “slashing” mechanism, in which a verifier is penalized for a constant amount when the verifier is detected to have passed the verification of an invalid proof, has a two-fold challenge as follows. The first one is that in a decentralized environment, there is not a trusted “root” that makes the decision, so the decision needs to be performed in a voting protocol, but it is hard to ensure the honesty of voters with theoretical guarantees; the second one is the Verifier’s Dilemma [8]:
Verifier’s Dilemma:
  • If a mechanism is (incentive-)secure against strategic provers, then no (rational) provers would cheat.
  • If no prover would cheat and the verification has a non-zero computational cost, then the verifier’s optimal strategy is to lazily report “success” without actual verification.
  • If all verifiers are rational and would not actually verify, then the security properties no longer hold.
Formally, the Verifier’s Dilemma can be formulated as the following theorem. The proof is deferred to Appendix A.
Theorem 1 (Verifier’s Dilemma).
In a verification game in which
  • A verifier’s report is binary, e.g. “Success” or “Fail”;
  • The verification result of an honest proof is always “Success”;
  • Honest verification has a strictly positive cost,
It is impossible to design a verification mechanism with a pure-strategy Nash equilibrium that the prover and verifier(s) simultaneously act honestly.
The Verifier’s Dilemma only occurs in the scenario of $\epsilon$, the probability of cheating, tends to zero, so that any bounded reward for catching a cheat cannot uniformly compensate for the verification cost in expectation. Hence, a conventional penalty-based approach can only lower, but not eliminate, the rate of cheating in the pool of proofs. For example, the study of [7] shows that the opML mechanism achieves a mixed-strategy Nash equilibrium that the prover has a constant probability to cheat and the verifier has a constant probability to honestly verify.
To achieve an honest pure-strategy Nash equilibrium, a natural idea is to introduce extra information or deliberate errors, e.g. insert deliberate objects (which can be valid or invalid), as so-called “flags” to incentivize verifiers to find and report, as in the works of [9, 10, 8, 11]. The simple flag-insertion mechanism works well for “honest-but-lazy” verifiers, as long as the expected reward of finding the flags can cover the verification cost. However, in the scenario where the validity of the verification result is also costly to verify, e.g., Proof-of-Learning [5, 11], we either need a sort of authentication or higher-level verifiers to verify their validity, or need additional phases of “committee voting” to decide on their validity. As long as we want to design a fully decentralized blockchain with no trusted “root” parties, there is no way to access the “ground truth” and the validity and payments can only be decided by consensus among the voting parties.
In the field of computational economics, the concept of peer prediction [12] refers to a wide scope of mechanisms to elicit honest information when direct verification is unavailable, which is widely adopted in the applications of dataset acquisition [13], peer grading [14], etc. A widely used practice for peer prediction is to design a mutual-information-based payment rule [15] and leverage the data processing inequality to make sure that any manipulation cannot increase the expected utility for an agent. While the logarithmic scoring rule in [13] is convenient for usage in scenarios with known prior, this family of peer prediction mechanisms is still unable to directly resolve the Verifier’s Dilemma because the infinitesimal cheating probability will make the logarithmic scoring rule diverge. In the practice of blockchain systems, as the Verifier’s Dilemma leads to an arbitrarily low cheating probability, its empirical value becomes difficult to estimate and the scoring rule’s sensitivity to small probabilities will severely undermine the robustness of the verification mechanism.
In contrast to most peer prediction mechanisms that guarantee Bayesian Nash equilibria requiring a known prior, a recent work [16] utilizes determinant-based mutual information (DMI) to design a dominantly-truthful finite-task peer prediction mechanism which has potential usage to resolve the Verifier’s Dilemma. However, although it does not need an estimation of $\epsilon$, its need and assumption of multiple homogeneous tasks, i.e. at least 4 tasks sharing a common $\epsilon$, may still lead to complication in the analysis as it has assumptions on and is still sensitive to the latent value of $\epsilon$. While we acknowledge the potential of the DMI mechanism as a possible solution to the problem, we would still be motivated to design a conceptually “more robust” mechanism that is insensitive to $\epsilon$ as long as it is small enough.
Compared to traditional peer prediction mechanisms which mainly consider incentive compatibility (IC), in our setting, we are additionally required to explicitly consider individual rationality (IR) so that an honest verifier would get an expected reward no less than the verification cost, and the verification cost may be dependent on the verification result. On the other hand, due to the permissionless property of the blockchain system, we also need a no-free-lunch (NFL) property that any verifier who does not actually perform the verification cannot get a positive expected reward from the mechanism.
In this research, we develop a theoretical framework with modeling of decentralized verification game (DVG), and initiate the study that combines the ideas of flags and peer prediction into our proposed mechanism, named capture-the-flag [^1] The term “Capture-The-Flag (CTF)” refers to a cybersecurity competition in which participants attempt to find “flags” hidden in deliberately vulnerable systems to claim rewards, which is conceptually similar to our design. peer prediction (CTF-PP), which only needs one phase in its procedure and simultaneously satisfies the following properties:
  • Interim [^2] In the DVG, a verifier’s action has two steps: first deciding on whether to verify, and then reporting a type. The term “interim” refers to considering the expected utility based on all the information she has based on her first-step decision. unique incentive compatibility (interim UniIC): Given all other verifiers act honestly, a verifier, after performing the verification, maximizes her expected utility when she reports honestly. Furthermore, if she reports a different type from her observation, her expected utility is non-positive.
  • Interim individual rationality (interim IR): Given all other verifiers act honestly, a verifier, after performing the verification, gets a non-negative expected utility when she acts honestly.
  • Interim no-free-lunch (interim NFL): Given all other verifiers act honestly, a verifier cannot get a positive expected utility without doing the verification.
As a preliminary result, we consider a simple case of 2-verifier DVG, and model the design of a desirable mechanism as a linear program. By computing a numerical solution to a typical set of parameters that applies to the model in the Incentive-Secure PoL [11], we show the potential of our approach as a solution that breaks the critical barrier of Verifier’s Dilemma in the design of blockchain (and more generally, decentralized) verification mechanisms.
In the rest of this paper, we omit the term “interim” as all the properties are considered in the interim setting.
2 Modeling: 2-Verifier Decentralized Verification Game
To initiate the study, we first consider a simple case with 2 verifiers. In a 2-verifier DVG, the verifiers $i=1,2$ independently verifies an on-chain “proof”, which has an underlying type $\theta\in S$ that may be “Honest” (00), “Flag $j$” ($F_{j}$, $j=1,2,\cdots,m$) or “Dishonest” ($1$). [^3] As discussed in the design of [11], the existence of at least two different flags can be necessary for detecting cheats. This statement may differ in different settings, but as a general model it is harmless to allow multiple flags. Since the system can insert flags to maintain a pre-set flag rate that robustly incentivizes verification, we have the prior probabilities $P(\theta=F_{i})$ as publicly known information; on the other hand, there is a small but unknown probability $\epsilon\in[0,\epsilon_{0}]$ that the proof is dishonest, i.e. $\Pr[\theta=1]=\epsilon$, with a known upper bound $\epsilon_{0}$. Hence, we have $\Pr[\theta=0]=1-\sum P(\theta=F_{i})-\epsilon$.
Similar to [11], we first consider the verification process with a lossy-channel model as follows. When verifier $i$ verifies the proof, the distribution of the observation $X_{i}$ is dependent on $\theta$ in this way:
  • A honest proof is always observed as honest, i.e. $\Pr[X_{i}=0|\theta=0]=1$.
  • A flag $F_{j}$ can be detected with known probability $\mu_{j}$ or missed and observed as honest, but will never be observed as dishonest or the other flag, i.e. $P(X_{i}=F_{j}|\theta=F_{j})=\mu_{j}$ and $P(X_{i}=0|\theta=F_{j})=1-\mu_{j}$.
  • A dishonest proof can be adversarial and observed as any type, but the probabilities are known to the public, and it can be correctly detected with a probability at least $\kappa>0$, i.e. $P(X_{i}=1|\theta=1)\geq\kappa$.
For each verifier $i$, she first makes a decision to either honestly verify the proof and get the observation, which gains her the access of $X_{i}$ but incurs a publicly-known cost $c(X_{i})$ which can depend on $X_{i}$, or does not verify and has no access to $X_{i}$. Hence, $i$’s belief on $Z_{-i}$ is the conditional distribution of $X_{-i}|X_{i}$ if she performs the verification, or the marginal distribution of $X_{-i}$ if she does not.
Then, she reports a $Z_{i}$ that maximizes $E_{\tilde{Z}_{-i}\sim\mathcal{B}}[R_{i}(Z_{i},\tilde{Z}_{-i})]$ in which $\mathcal{B}$ is her belief of $Z_{-i}$, and claims that $Z_{i}$ is her observation. After each verifier $i$ independently reports $Z_{i}$ without seeing $Z_{-i}$ [^4] This can be implemented via a cryptographic commitment scheme., the system has the information of $Z_{1},Z_{2}$, but not $\theta$, and rewards each prover $Z_{i}$ based on a scoring rule from the reports, denoted as $R_{i}(Z_{i},Z_{-i})$. Then, verifier $i$’s net utility is $R_{i}(Z_{i},Z_{-i})-c(X_{i})$ if she honestly verified, or $R_{i}(Z_{i},Z_{-i})$ if she did not actually verify. In the rest of this section, we consider the two cases separately.
2.1 IR and UniIC Constraints for Non-Lazy Verifiers
The IR constraint requires that given the verifier $i$ observes $X_{i}$, truthfully reporting it gains her an expected reward no less than $c(X_{i})$. Since $X_{i}$ and $X_{-i}$ are independent conditioned on $\theta$, Define $r_{X_{i}}(Z_{i})$ as verifier $i$’s expected reward of reporting $Z_{i}$ conditioned on observing $X_{i}$, then $r_{X_{i}}(Z_{i})$ can be computed as
$r_{X_{i}}(Z_{i})$ $=\sum_{X_{-i}\in S}R_{i}(Z_{i},X_{-i})P(X_{-i}|X_{i})$ (1)
$=\sum_{X_{-i}\in S}R_{i}(Z_{i},X_{-i})\frac{P(X_{i},X_{-i})}{P(X_{i})}$ (2)
$=\sum_{X_{-i}\in S}R_{i}(Z_{i},X_{-i})\frac{\sum_{\theta\in S}P(\theta)P(X_{i}|\theta)P(X_{-i}|\theta)}{\sum_{\theta\in S}P(\theta)P(X_{i}|\theta)}.$ (3)
While the probabilities are dependent on $\epsilon$, as long as $\epsilon$ is small enough, for $X_{i}\in\{0,F_{1},F_{2}\}$, the $r_{X_{i}}$ is a continuous function w.r.t. $\epsilon$, so the IR constraints on $X_{i}\in\{0,F_{1},F_{2}\}$ can be implied by
$\sum_{X_{-i}\in S}R_{i}(X_{i},X_{-i})\frac{\sum_{\theta\in S}\tilde{P}(\theta)\tilde{P}(X_{i}|\theta)\tilde{P}(X_{-i}|\theta)}{\sum_{\theta\in S}\tilde{P}(\theta)\tilde{P}(X_{i}|\theta)}\geq c(X_{i})+\delta,\qquad\forall X_{i}\in\{0,F_{1},F_{2}\},$ (4)
in which $\tilde{P}$ denotes the probability assuming $\epsilon=0$, and for any $\delta>0$, the constraints hold for some $\epsilon_{0}>0$. The condition is also necessary when $\delta=0$. In the rest of this paper, we say a condition is “sufficient and almost necessary” when it is sufficient with a $\epsilon_{0}>0$ depending on $\delta$, and when we set $\delta=0$, it becomes a necessary condition.
For the case of $X_{i}=1$, from the DVG model, we know that it must hold that $\theta=1$. Therefore, we have
$r_{1}(1)$ $=\sum_{X_{-i}\in S}R_{i}(1,X_{-i})P(X_{-i}|\theta=1)$ (5)
$\geq c(1)+\delta.$ (6)
So Eq. (4) and Eq. (5) are sufficient and almost necessary conditions that a CTF-PP mechanism is IR.
For the IC constraint, we need and only need $r_{X_{i}}(X_{i})=\max_{Z_{i}\in S}\{r_{X_{i}}(X_{i})\}$. With similar arguments, we can also develop sufficient and almost necessary conditions that a CTF-PP mechanism is IC. Actually, given that the IR is satisfied we can define a stronger notion of Uniquely-IC as follows:
  • Uniquely IC (UniIC): In addition to the IC requirement, given all other verifiers act honestly and a verifier has already performed the verification, then she gets a negative expected utility when she reports any type different from her observation.
Besides conventional IC notions, the UniIC requirement additionally rules out the possibility that a dishonest verifier cheats the system without losing money. Assuming that the IR constraints are already satisfied, the UniIC constraints can be formulated as:
$r_{X_{i}}(Z_{i})\leq c(X_{i})-\delta,\qquad\forall X_{i}\in S,\>Z_{i}\neq X_{i}.$ (7)
2.2 NFL Constraints for Lazy Verifiers
We assume verifier $-i$ is honest, i.e. $Z_{-i}=X_{-i}$. If verifier $i$ does not actually verify the proof, she has no information on $\theta$ and her strategy can only be reporting any type in $S=\{0,F_{1},\cdots,F_{m},1\}$, or any convex combination of them. Hence, $i$’s utility when she lazily reports $Z_{i}$ is denoted as:
$r_{\bot}(Z_{i})=\sum_{X_{-i}\in S}R_{i}(Z_{i},X_{-i})P(X_{-i}).$ (8)
From the NFL requirement and assuming small $\epsilon\leq\epsilon_{0}$, a sufficient and almost necessary condition is the following linear constraints
$\sum_{X_{-i}\in S}R_{i}(Z_{i},X_{-i})P(X_{-i})\leq-\delta,\qquad\forall Z_{i}\in S.$ (9)
3 Theoretical Guarantee for 2-Verifier DVG: LP Modeling and Feasibility
In this section, we show a preliminary result that the desirable CTF-PP mechanism exists for any 2-verifier DVG that satisfies mild conditions.
[…]
We define decision variable as the scoring matrix $T:S^{2}\to\mathbb{R}$ with $T_{xy}=R_{i}(x,y)$, and denote $W=BT^{\prime}$. Then $W_{xy}=r_{x}(y)$, which is the expected reward verifier $i$ gets from the mechanism when she observes $x$ and reports $y$. The IR and UniIC conditions are equivalent to the following:
$W_{xx}$ $\geq c(x),\qquad\forall x\in S;$ (10)
$W_{xy}$ $\leq c(x),\qquad\forall x\in S,\quad y\in S-\{x\}.$ (11)
Similarly, we denote $W_{\bot}=B_{\bot}T^{\prime}$, then $W_{\bot y}=r_{\bot}(y)$ is the expected reward verifier $i$ gets when she does not verify and lazily reports $y$. Then the NFL conditions are equivalent to the following:
$W_{\bot}\leq 0.$ (12)
Hence, we only need to find a feasible solution of the linear system (10)-(12), with a small margin $\delta$ to each constraint. Here, we propose our main theorem, with the proof deferred to Appendix B.:
Theorem 2 (Main Theorem).
If $B$ is invertible, and $P(X_{i}=y|\theta=1)=0$ for any $y\neq 1$ (i.e. a no-cheat proof is never observed as a cheat), then for some $\epsilon_{0}>0$, there exists a fixed CTF-PP mechanism for the 2-verifier DVG that is IR, NFL and UniIC for any $\epsilon\in[0,\epsilon_{0}]$.
Particularly, we show that our method always works for the DVG with the lossy-channel model as defined in Section 2 with the following proposition. The proof is omitted.
Proposition 1.
In the lossy-channel model defined in Section 2, the cheat-free belief matrix $B$ is invertible.
$X_{i}=0$ $X_{i}=F_{1}$ $X_{i}=F_{2}$ $X_{i}=1$
$\theta=0$ $1$ 00 00 00
$\theta=F_{1}$ $\frac{1}{2}$ $\frac{1}{2}$ 00 00
$\theta=F_{2}$ $\frac{1}{2}$ 00 $\frac{1}{2}$ 00
$\theta=1$ $\frac{1}{2}$ $\frac{1}{8}$ $\frac{1}{8}$ $\frac{1}{4}$
Table 1: $P(X_{i}|\theta)$, the observation distribution conditioned on $\theta$.
4 Experimental Results: Verification of Incentive-Secure PoL
In this section, we empirically demonstrate the process of designing a CTF-PP mechanism, for one set of parameters that is useful for practical interest.
4.1 Construction of the Scoring Rule
$X_{2}=0$ $X_{2}=F_{1}$ $X_{2}=F_{2}$ $X_{2}=1$
$X_{1}=0$ $\frac{5}{8}$ $\frac{1}{16}$ $\frac{1}{16}$ 00
$X_{1}=F_{1}$ $\frac{1}{16}$ $\frac{1}{16}$ 00 00
$X_{1}=F_{2}$ $\frac{1}{16}$ 00 $\frac{1}{16}$ 00
$X_{1}=1$ 00 00 00 00
Table 2: Joint probabilities $P(X_{1},X_{2})$ ($\epsilon=0$).
$X_{2}=0$ $X_{2}=F_{1}$ $X_{2}=F_{2}$ $X_{2}=1$
$X_{1}=0$ $\frac{5}{6}$ $\frac{1}{12}$ $\frac{1}{12}$ 00
$X_{1}=F_{1}$ $\frac{1}{2}$ $\frac{1}{2}$ 00 00
$X_{1}=F_{2}$ $\frac{1}{2}$ 00 $\frac{1}{2}$ 00
$X_{1}=1$ $\frac{1}{2}$ $\frac{1}{8}$ $\frac{1}{8}$ $\frac{1}{4}$
Table 3: The cheat-free belief matrix $B$.
We consider the 2-verifier DVG which captures the case of one stage in [11]. Here, we set the distribution $\theta$ as $P(\theta=F_{1})=P(\theta=F_{2})=\frac{1}{4}$, which means that half of all stages are flagged. Then we consider the lossy-channel model in which $\mu_{1}=\mu_{2}=\frac{1}{2}$, as each verifier independently chooses half of all stages  [^6] It is significantly more than needed, but does work. and each flag is detected with probability $1$ when verified. According to the CTF protocol, when a cheating stage is chosen by a verifier, it has an $\frac{1}{2}$ chance to be correctly detected, and a $\frac{1}{4}$ chance to be observed as $F_{1}$ or $F_{2}$ respectively. Hence, $P(X_{i}|\theta)$ is shown as in Table 1, and assuming $\epsilon=0$, we compute the marginal distribution of $X_{i}$ as $B_{\bot}=[\frac{3}{4},\frac{1}{8},\frac{1}{8},0]$.
Then, assuming $\epsilon=0$, from $P(X_{i},X_{-i})=\sum_{\theta}P(\theta)P(X_{i}|\theta)P(X_{-i}|\theta)$ we can compute the joint probabilities $P(X_{1},X_{2})$, as shown in Table 2.We accordingly compute the post-observation belief $P(X_{2}|X_{1})=\frac{P(X_{1},X_{2})}{P(X_{1})}$ for $X_{1}\neq 1$, and $P(X_{2}|X_{1}=1)=P(X_{2}|\theta=1)$, getting the cheat-free belief matrix $B$ as Table 3.
From the nature of the CTF mechanism, in which the observation of $F_{1},F_{2}$ of $1$ takes twice the computational cost of a stage, we set $c(F_{1})=c(F_{2})=c(1)=2c$. On the other hand, the “observation” of a 00 has two cases: one is that the verifier has verified the stage that is not cheated or flagged, which has a $\mu(1-\eta)=\frac{1}{4}$ probability, and one is that the verifier does not verify the stage from the random verification protocol, which has a $1-\mu=\frac{1}{2}$ probability. Hence, we have an “amortized” $c(0)=\frac{1}{3}c$. Without loss of generality, we set $c=1$.
With the knowledge of $B,B_{\bot}$ and $c$, Eqs. (10)-(12) are the constraints that a desirable scoring rule for a CTF-PP mechanism should satisfy. For the robustness of our mechanism, we do not want the payments to have extremely large absolute values. Hence, we construct the linear program as:
$\qquad\qquad minimize~{}\qquad~{}$ $M$
$\qquad\qquad s.t.~{}\qquad~{}$ (10)-(12)$,\qquad-M\leq T\leq M.( )-$
We set a margin of $\delta=0.2$, and compute a numerical solution to this LP, getting a scoring rule as shown in Table 4.
$Z_{-i}=0$ $Z_{-i}=F_{1}$ $Z_{-i}=F_{2}$ $Z_{-i}=1$
$Z_{i}=0$ $+2.0959$ $-7.1643$ $-7.1643$ $-1.0764$
$Z_{i}=F_{1}$ $-1.5421$ $+6.4732$ $-4.4574$ $-1.2378$
$Z_{i}=F_{2}$ $-1.5421$ $-4.4574$ $+6.4732$ $-1.2378$
$Z_{i}=1$ $-2.2000$ $+5.8000$ $+5.8000$ $+7.4000$
Table 4: $T_{Z_{i}Z_{-i}}=R_{i}(Z_{i},Z_{-i})$ as a numerical solution.
4.2 Evaluation
With this scoring rule, given verifier $-i$ acts honestly, we report the expected utility of verifier $i$ is in Table 5, assuming $\epsilon=0$, showing that the verifier gets a positive expected utility if and only if she verifies and reports honestly.
Furthermore, we consider the case $\epsilon>0$. We plot the maximum expected utility of dishonest actions and the minimum expected utility of honest actions in Figure 1. From the plot, we show that the introduction of the margin keeps the IR, UniIC and NFL properties of our mechanism as long as $\epsilon<0.045$. Hence, we demonstrate that our design of the CTF-PP mechanism for the 2-verifier DVG can incentivize honest verification even if there is no dishonest prover, thus bypassing the Verifier’s Dilemma and achieving a pure-strategy Nash equilibrium that the prover and verifiers simultaneously act honestly.
Reporting 00 Reporting $F_{1}$ Reporting $F_{2}$ Reporting $1$
Observing 00 $\mathbf{+0.2192}$ $-1.4504$ $-1.4504$ $-1.2000$
Observing $F_{1}$ $-4.5342$ $\mathbf{+0.4655}$ $-4.9998$ $-0.2000$
Observing $F_{2}$ $-4.5342$ $-4.9998$ $\mathbf{+0.4655}$ $-0.2000$
Observing $1$ $-3.0122$ $-2.8285$ $-2.8285$ $\mathbf{+0.2000}$
Does not verify $-0.2192$ $-0.9046$ $-0.9046$ $-0.2000$
Table 5: Verifier’s expected utility, $\epsilon=0$.
Figure 1: Verifier’s expected utility, $\epsilon>0$.
5 Discussion
In this paper, we develop a theoretical framework for the decentralized verification game on blockchain proving protocols and get preliminary results that the combination of the Capture-The-Flag design (as in [11]) and peer prediction techniques has the potential to resolve the Verifier’s Dilemma in a fully decentralized environment. On the other hand, we also explore the design of peer prediction mechanisms with broader agent strategy spaces and more general settings (e.g. observation costs dependent on observation results). In future work, we will improve and broaden the study in the following aspects:
  1. 1.
    Generalize our setting to multiple ($\geq 3$) verifiers and a wider scope of noisy-observation models, e.g., the observation distribution of a cheat is unknown (but still with a $\geq\kappa$ probability to be correctly detected).
  2. 2.
    While this paper is mainly on the elicitation of truthful verification results, we will also develop voting/aggregation mechanisms that (optimally) make decisions on whether to accept the proof.
  3. 3.
    From the anonymous nature of the blockchain systems, it is also interesting to explore collusion-proof mechanisms for DVGs and (general) peer prediction, similar to the prominent line of research on collusion-proof transaction fee mechanisms (TFMs) [17, 18, 19].
Additionally, while our mechanism is conceptually designed for $\epsilon\approx 0$ and does not work for a relatively large $\epsilon$ (e.g. $>0.045$ in the example), in real-world cases where $\epsilon$ is (unexpectedly) large but estimable, the system can be augmented to adjust the belief matrix $B$ based on an (automated) estimation of $\epsilon$, denoted as $\hat{\epsilon}$, and compute an adjusted scoring rule accordingly. From the same argument on the margin $\delta$, the adjusted scoring rule would incentivize honest verification as long as $|\epsilon-\hat{\epsilon}|$ is small.
Acknowledgments
Zishuo Zhao would like to thank Yuqing Kong, Shuran Zheng and Hongyin Chen for insightful discussions on the design of peer prediction mechanisms, and Zhixuan Fang for discussions on topics of decentralized consensus in blockchain systems.
Appendix
Appendix A Proof of Theorem 1
Assume we have such a mechanism. By the definition of Nash equilibrium, we consider a fixed verifier. Given that the prover and all other verifiers act honestly, that verifier should be incentivized to do the honest verification.
Since the prover is honest, when that verifier performs honest verification, the result should always be “Success”. However, suppose the verifier simply reports “Success” without verification. In that case, the outcome is the same but the verifier saves the verification cost, so the verifier is incentivized to deviate from the honest strategy.
That leads to a contradiction, so no such mechanism exists.
Appendix B Proof of Theorem 2
Notice that if $B$ is invertible, for any given $W:S^{2}\to\mathbb{R}$, we can compute a $T=(B^{-1}W)^{\prime}$ that satisfies $W=BT^{\prime}$. Since we have
$B_{\bot y}$ $=P(X_{-i}=y)$ (13)
$=\sum_{x\in S}P(X_{i}=x)\cdot P(X_{-i}=y|X_{i}=x)$ (14)
$=\sum_{x\in S}B_{\bot x}\cdot B_{xy},$ (15)
It holds that $B_{\bot}=B_{\bot}B$. Hence, $W_{\bot}=B_{\bot}T^{\prime}=B_{\bot}BT^{\prime}=B_{\bot}W$. Since $B_{\bot}\geq 0$, $W_{\bot}$ is a convex combination of rows in $W$. Therefore, we only need to construct a $W$ that satisfies Eq. 10 with non-diagonal entries small enough.
Here, for a constant $M>0$ large enough, we construct $W$ as:
$W_{xx}$ $=c(x)+\delta,\qquad\forall x\in S;$ (16)
$W_{xy}$ $=-M,\qquad\quad~{}~{}\,\forall x\in S,\quad y\in S-\{x\}.$ (17)
Then,
$W_{\bot y}$ $=\sum_{x\in S}B_{\bot x}W_{xy}$ (18)
$=B_{\bot y}W_{yy}+\sum_{x\in S-\{y\}}B_{\bot x}W_{xy}$ (19)
$=B_{\bot y}(c(y)+\delta)-(1-B_{\bot y})M.$ (20)
Since the existence of flags introduces randomness in the observation, we have $\max\{B_{\bot}\}<1$. Hence, denote $B^{*}=\max\{B_{\bot y}\}$, we only need to let
$M\geq\frac{B^{*}\cdot(c(y)+\delta)+\delta}{1-B^{*}},$ (21)
Then the required constraints of Eqs. (10)-(12) are satisfied with a margin of $\delta$.
Then we consider the scenario that $\epsilon>0$ but is small enough. In this case, define $B(\epsilon)$ and $B_{\bot}(\epsilon)$ as the belief matrix and blind-belief matrix considering the influence of $\epsilon$. We can see that for any $x\neq 1$, since $\tilde{P}(X_{i}=x)>0$, the influences of $\epsilon$ on $P(X_{-i}|X_{i}=x)$ and $P(X_{-i})$ are upper bounded by $O(\epsilon)$, and because $\theta\neq 1\implies X_{i}\neq 1$, Eq. (5) always holds for any $\epsilon$. Therefore, the margin of $\delta$ ensures that the constraints are not violated as long as $\epsilon_{0}$ is small enough.
Finally, let $T=(B^{-1}W)^{\prime}$, then $T$ is a scoring rule that satisfies the requirements.